# Global Continuation of Stable Periodic Orbits in Systems of Competing Predators

### v1

@@ 3 Proof of stability
We assume that Corollary 2.6 holds such that there exists a locally unique $\tilde{\chi}=(\tilde{\tau},\tilde{\zeta}_{1},\tilde{\zeta}_{2},\tilde{u})\in\mathcal{X}_{\nu}$ satisfying $F(\tilde{\chi})=0$. In this section, we show how to solve the eigenproblem for the whole branch of periodic orbits, and provide a sufficient criterion for such a family to be comprised of stable periodic orbits. The underlying argument of the proof is similar to the one used for the proof of existence in Section 2, in that it hinges on a local contraction argument. However, the boundary crossing (where transcritical bifurcations occur) requires us to verify stability by different means when the branch is “close to” and “far from” the boundary; the adjectives “close” and “far” are to be quantified in this section.
@@
$\mathcal{Y}_{\nu}:=\textnormal{Mat}_{3\times 3}(\mathcal{P}_{\nu})\times\textnormal{Mat}_{3\times 3}(\mathcal{U}_{\nu}),$ (73)
@@
$\|\upsilon\|_{\mathcal{Y}_{\nu}}:=\sum_{i,j=1}^{3}\|C_{i,j}\|_{\nu}+\sum_{i=1}^{3}\sum_{j=1}^{3}\|V_{i,j}\|_{\mathcal{U}_{\nu}},\qquad\text{for all }\upsilon=(C,V)\in\mathcal{Y}_{\nu}.$ (74)
@@
$G(C,V):=\begin{pmatrix}\Pi_{N,K}V(0)-I_{3\times 3}\\ \partial_{t}V+VC-\tilde{\tau}\partial_{u}f(\tilde{\zeta}_{1},\tilde{\zeta}_{2},\tilde{u})V\end{pmatrix}.$ (75)
If $G(\tilde{C},\tilde{V})=0$, then we can control the spectrum of $\tilde{C}$ by using the Gershgorin circle theorem. However, as we cross the boundary planes $Q_{1},Q_{2}$ at respectively $\hat{\eta}_{1},\hat{\eta}_{2}$, one eigenvalue crosses the imaginary axis through the origin $0\in\mathbb{C}$ due to the transcritical bifurcation. Thus, in a neighbourhood of $\hat{\eta}_{j}$, the sign of the real part of this “crossing eigenvalue” is undetermined. Recall that in Section 2.4 we have considered a slightly smaller interval $[h_{1}^{-},h_{2}^{+}]\subset[-1,1]$ which we partitioned into the three pieces:
@@
    $[h_{1}^{-},h_{1}^{+}]$ where the family crosses $Q_{1}$ at $\hat{\eta}_{1}\in(h_{1}^{-},h_{1}^{+})$;
@@
    $[h_{1}^{+},h_{2}^{-}]$ where the family is in the interior +$\mathbb{R}^{3}_{+}$;
@@
    $[h_{2}^{-},h_{2}^{+}]$ where the family crosses $Q_{2}$ at $\hat{\eta}_{2}\in(h_{2}^{-},h_{2}^{+})$.
We show that for all $\eta\in[h_{1}^{+},h_{2}^{-}]$ the family of periodic orbits is stable, and in $[h_{1}^{-},h_{1}^{+}]$ and $[h_{2}^{-},h^{+}_{2}]$ we verify that one eigenvalue remains in a bounded region of the left-half plane of $\mathbb{C}$. The procedure consists in the following steps.
@@
    We solve (72) for all $\eta\in[-1,1]$ by finding a zero $(\tilde{C},\tilde{V})\in\mathcal{Y}_{\nu}$ of $G$. To prove the existence of a zero of $G$, we rely on a contraction argument in the vein of Section 2; in fact, since $G$ is quadratic, the assumptions of Lemma 2.4 and Lemma 2.5 simplify, and we postpone the details to Appendix A.
@@
    We repeat the strategy of Step 2 on $[\eta_{1}^{-},\eta_{1}^{+}]$ (resp., $[\eta_{2}^{-},\eta_{2}^{+}]$) to show that one eigenvalue of $\tilde{C}(\eta)$ remains in a compact region of $\{z\in\mathbb{C}\,:\,\mathrm{Re}(z)<0\}$ for all $\eta\in[\eta_{1}^{-},\eta_{1}^{+}]$ (resp., $\eta\in[\eta_{2}^{-},\eta_{2}^{+}]$). This suffices to show that the family of periodic orbits loses its stability exactly at the crossing of $Q_{1}$ and $Q_{2}$ occurring at $\hat{\eta}_{1}$ and $\hat{\eta}_{2}$ respectively. Indeed, only one eigenvalue can cross the imaginary axis. Since this occurs through the origin 00 at $Q_{1}$ and $Q_{2}$, any additional crossing of the imaginary axis must also pass through the origin 00. The contraction argument guarantees that $DF(\tilde{\chi})$ is invertible which, in particular, forbids the two non-trivial Floquet exponents $\mu_{1}$ and $\mu_{2}$ to vanish except on $Q_{1}$ and $Q_{2}$ – each being crossed exactly once.
@@
To establish the stability of the family of periodic orbits described in Theorem 1.1, we used a Fourier–Chebyshev approximation $\bar{\upsilon}=(\bar{C},\bar{V})\in\Pi_{N,K}\mathcal{Y}_{\nu}$ with $K=20$, $N=30$. The exact solution $\tilde{\upsilon}=(\tilde{C},\tilde{V})$ to (72) lies within a distance $\|\tilde{\upsilon}-\bar{\upsilon}\|_{C^{0}}\leq\|\tilde{\upsilon}-\bar{\upsilon}\|_{\mathcal{Y}_{\nu}}\leq 5\times 10^{-5}$; see [33]. Figure 2 shows the non-trivial eigenvalues of $\tilde{C}$, i.e. the Floquet exponents $\mu_{1}$ and $\mu_{2}$.
@@ 4 Outlook and future work
Theorem 1.1 provides a global family of positive periodic orbits in the shape of a tube, whose stable periodic orbits reach a boundary limit cycle; see Figure 1. Nevertheless, through (typically multi-parameter) continuations, other invariant subsets or more intricate dynamics near our global family could possibly occur. As main examples, periodic orbits may undergo bifurcations such as a Neimark–Sacker, a period-doubling, a homoclinic bifurcation, or a blue sky catastrophe. See for instance [38, Chapter 5–7] and the references therein. As far as we know, there is no evidence of Neimark–Sacker bifurcation giving rise to invariant tori, or blue-sky catastrophes; but there is evidence of a cascade of period-doubling bifurcations; see [20, 37].
As a matter of fact, using both $\kappa$ and $a_{1}$ as adjustable parameters, the proven global family at $a_{1}=10$ (see Figure 1) seems to eventually lose its stability in the interior when $a_{1}$ decreases. We did not attempt to use our continuation method to rigorously track the value of $a_{1}$ at which the loss of stability occurs, as the global branch becomes increasingly challenging to approximate (see e.g., Figure 3(a)). Achieving a rigorous proof in this regime requires taking a substantially larger number of Fourier modes $K$ and Chebyshev modes $N$, thereby significantly increasing the computational cost of the continuation method developed in this article. Yet, numerical evidence suggests that this instability arises from the occurrence of period-doubling bifurcations along the branch; see Figure 3(b). It may be the case that further decreasing $a_{1}$ results in a cascade of period-doubling bifurcations, eventually leading to a chaotic attractor in $\mathbb{R}_{+}^{3}$ around $a_{1}\approx 4$; see Figure 3(c). If there is indeed a chaotic attractor, its existence, and the mechanism behind its birth deserve further study.
@@
Figure 3: The parameter values are $a_{2}=41$, $d_{1}=0.8$, $d_{2}=0.5$, and $m_{1}=m_{2}=y_{1}=y_{2}=\gamma=1$. (a) Numerical approximation of the global family for $a_{1}=6$. The orange rings corresponds to a numerical observation of a period-doubling bifurcation. (b) Numerical approximation of the Floquet multipliers associated to (a). The orange cross marks the crossing through $-1$, i.e., a potential period-doubling bifurcation. (c) Projection onto the $(X_{1},X_{2})$-plane of a numerical chaotic attractor in +$\mathbb{R}^{3}_{+}$ for $a_{1}=4$.
@@ 5 Acknowledgments
We are grateful to Sze-Bi Hsu for the suggestion of solving the stable connection problem and to Bernold Fiedler for many inspiring discussions. J.-Y. Dai was supported by NSTC (National Science and Technology Council) grant 113-2628-M-007-005-MY4. O. Hénot was supported by the Fondation Mathématique Jacques Hadamard and the ANR project CAPPS: ANR-23-CE40-0004-01. P. Lappicy was supported by Marie Skłodowska–Curie Actions, UNA4CAREER H2020 Cofund, 847635, with the project DYNCOSMOS. N. Vassena was supported by the DFG (German Research Society), project No. 512355535. O. Hénot also thank the National Center for Theoretical Sciences for its support during the preparation of this work.
@@ Appendix A Contraction argument for the Floquet normal form
In the context of this manuscript, we want to verify Lemma 2.4 and Lemma 2.5 in the Banach space $\mathcal{Y}_{\nu}$ and the fixed-point operator
@@
thereby obtaining a zero of the mapping $G$ defined in 75 via Corollary 2.6. Similarly to the construction of the injective linear operator $A$, we set $B\in\mathscr{B}(\mathcal{Y}_{\nu},\mathcal{Y}_{\nu})$ as
$B:=B_{\textnormal{finite}}\Pi_{K}+B_{\textnormal{tail}}\Pi_{>K},$ (78)
@@
$B_{\textnormal{tail}}\upsilon:=(0,I_{3\times 3}\otimes V^{\prime}),\qquad(V_{ij}^{\prime})_{k}:=\begin{cases}0,&|k|\leq K,\\ (ik)^{-1}(V_{ij})_{k},&|k|>K,\end{cases}\quad i,j=1,2,3.$ (79)
@@
$DG(\upsilon)=\begin{pmatrix}0&I_{3\times 3}\otimes E_{0}\Pi_{N,K}\\ I_{3\times 3}\otimes V&I_{3\times 3}\otimes\partial_{t}+C^{T}\otimes I_{3\times 3}-I_{3\times 3}\otimes\tilde{\tau}\partial_{u}f(\tilde{\zeta}_{1},\tilde{\zeta}_{2},\tilde{u})\end{pmatrix},$
where $C^{T}$ is the 3-by-3 matrix transpose of $C$ and $E_{0}:\textnormal{Mat}_{3\times 3}(\mathcal{U}_{\nu})\to\textnormal{Mat}_{3\times 3}(\mathcal{P}_{\nu})$ is the evaluation operator at 00, i.e. $E_{0}V:=V(0)$.
Formulas for $\|BG(\bar{\upsilon})\|_{\mathcal{Y}_{\nu}}$, $\|BDG(\bar{\upsilon})-I\|_{\mathscr{B}(\mathcal{Y}_{\nu},\mathcal{Y}_{\nu})}$, and $\sup_{\upsilon\in\textnormal{cl}(B_{R}(\bar{\upsilon}))}\|B(DG(\upsilon)-DG(\bar{\upsilon})\|_{\mathscr{B}(\mathcal{Y}_{\nu},\mathcal{Y}_{\nu})}$ are in fact much simpler than those obtained for $F$ in Sections 2.3.1, 2.3.2 and 2.3.3. Indeed, $G$ has only the single quadratic term $VC$. Thus, we can freely pick $R=\infty$, and $\|B(DG(\upsilon)-DG(\bar{\upsilon}))\|_{\mathscr{B}(\mathcal{Y}_{\nu},\mathcal{Y}_{\nu})}\leq 2\|B\|_{\mathscr{B}(\mathcal{Y}_{\nu},\mathcal{Y}_{\nu})}$ for all $\upsilon\in\mathcal{Y}_{\nu}$. Then, in the context of the fixed-point operator $\upsilon\mapsto\upsilon-BG(\upsilon)$, the equivalence of Lemma 2.4 and Lemma 2.5 read: finding $r\in(0,\infty)$, such that
$\|BG(\bar{\upsilon})\|_{\mathcal{Y}_{\nu}}+r\|BDG(\bar{\upsilon})-I\|_{\mathscr{B}(\mathcal{Y}_{\nu},\mathcal{Y}_{\nu})}+r^{2}\|B\|_{\mathscr{B}(\mathcal{Y}_{\nu},\mathcal{Y}_{\nu})}$ $\leq r,$
$\|BDG(\bar{\upsilon})-I\|_{\mathscr{B}(\mathcal{Y}_{\nu},\mathcal{Y}_{\nu})}+2r\|B\|_{\mathscr{B}(\mathcal{Y}_{\nu},\mathcal{Y}_{\nu})}$ $<1.$
@@
$\hat{W}_{1}:=\begin{pmatrix}0&I_{3\times 3}\otimes E_{0}\Pi_{N,K}\\ I_{3\times 3}\otimes\bar{V}&I_{3\times 3}\otimes\partial_{t}+\bar{C}^{T}\otimes I_{3\times 3}-\omega_{1}\end{pmatrix},$ (80)

### v2

@@ 3 Proof of stability
We assume that Theorem 2.1 holds such that there exists a locally unique $\tilde{\chi}=(\tilde{\tau},\tilde{\zeta}_{1},\tilde{\zeta}_{2},\tilde{u})\in\mathcal{X}_{\nu}$ satisfying $F(\tilde{\chi})=0$. In this section, we show how to solve the eigenproblem for the whole branch of periodic orbits, and provide a sufficient criterion for such a family to be comprised of stable periodic orbits. The underlying argument of the proof is similar to the one used for the proof of existence in Section 2, in that it hinges on a local contraction argument. However, the boundary crossing (where transcritical bifurcations occur) requires us to verify stability by different means when the branch is “close to” and “far from” the boundary; the adjectives “close” and “far” are to be quantified in this section.
@@
$\mathcal{Y}_{\nu}\coloneqq\textnormal{Mat}_{3\times 3}(\mathcal{P}_{\nu})\times\textnormal{Mat}_{3\times 3}(\mathcal{U}_{\nu}),$ (79)
@@
$\|\upsilon\|_{\mathcal{Y}_{\nu}}\coloneqq\sum_{i,j=1}^{3}\|C_{i,j}\|_{\nu}+\sum_{i=1}^{3}\sum_{j=1}^{3}\|V_{i,j}\|_{\mathcal{U}_{\nu}},\qquad\text{for all }\upsilon=(C,V)\in\mathcal{Y}_{\nu}.$ (80)
@@
$G(C,V)\coloneqq\begin{pmatrix}\Pi_{N,K}V(0)-I_{3\times 3}\\ \partial_{t}V+VC-\tilde{\tau}\partial_{u}f(\tilde{\zeta}_{1},\tilde{\zeta}_{2},\tilde{u})V\end{pmatrix}$. (81)
If $G(\tilde{C},\tilde{V})=0$, then we can control the spectrum of $\tilde{C}$ by using the Gershgorin circle theorem. However, as we cross the boundary planes $Q_{1},Q_{2}$ at respectively $\hat{\eta}_{1},\hat{\eta}_{2}$, one eigenvalue crosses the imaginary axis through the origin $0\in\mathbb{C}$ due to the transcritical bifurcation. Thus, in a neighbourhood of $\hat{\eta}_{j}$, the sign of the real part of this “crossing eigenvalue” is undetermined. Recall that in Section 2.3 we have considered a slightly smaller interval $h_{1}^{-},h_{2}^{+}]\subset[-1,1]$ which we partitioned into the three pieces:
@@
    $h_{1}^{-},h_{1}^{+}]$ where the family crosses $Q_{1}$ at $\hat{\eta}_{1}\in(h_{1}^{-},h_{1}^{+})$;
@@
    $h_{1}^{+},h_{2}^{-}]$ where the family is in the interior $\mathbb{R}^{3}_{+}$;
@@
    $h_{2}^{-},h_{2}^{+}]$ where the family crosses $Q_{2}$ at $\hat{\eta}_{2}\in(h_{2}^{-},h_{2}^{+})$.
We show that for all $\eta\in[h_{1}^{+},h_{2}^{-}]$ the family of periodic orbits is stable, and in $h_{1}^{-},h_{1}^{+}]$ and $h_{2}^{-},h^{+}_{2}$] we verify that one eigenvalue remains in a bounded region of the left-half plane of $\mathbb{C}$. The procedure consists in the following steps.
@@
    We solve (78) for all $\eta\in[-1,1]$ by finding a zero $(\tilde{C},\tilde{V})\in\mathcal{Y}_{\nu}$ of $G$. To prove the existence of a zero of $G$, we rely on a contraction argument in the vein of Section 2; in fact, since $G$ is quadratic, verifying the assumptions of Theorem 2.1 is easier, and we postpone the details to Appendix A.
@@
[…]
@@
[…]
@@ 4 Outlook and future work
Theorem 1.1 provides a global family of positive periodic orbits in the shape of a tube, whose stable periodic orbits reach a boundary limit cycle; see Figure 1. Nevertheless, through (typically multi-parameter) continuations, other invariant subsets or more intricate dynamics near our global family could possibly occur. As main examples, periodic orbits may undergo bifurcations such as a Neimark–Sacker, a period-doubling, a homoclinic bifurcation, or a blue sky catastrophe. See for instance [46, Chapter 5–7] and the references therein. As far as we know, there is no evidence of Neimark–Sacker bifurcation giving rise to invariant tori, or blue-sky catastrophes; but there is evidence of a cascade of period-doubling bifurcations; see [24, 45].
As a matter of fact, using both $\kappa$ and $a_{1}$ as adjustable parameters, the proven global family at $a_{1}=10$ (see Figure 1) seems to eventually lose its stability in the interior when $a_{1}$ decreases. We did not attempt to use our continuation method to rigorously track the value of $a_{1}$ at which the loss of stability occurs, as the global branch becomes increasingly challenging to approximate.
As a visual cue, Figure 3(a) shows that the periodic orbits transition from $Q_{1}$ to $Q_{2}$ quite abruptly. This discrepancy between the two boundary planes “kinks” the tube of solutions, as the Fourier–Chebyshev coefficients of the solution decay much slower. Thus, achieving a proof in this regime requires a substantially larger number of Fourier modes $K$ and Chebyshev modes $N$. While the use of fast Fourier transform algorithms mitigate the computational complexity with respect to the parameter (essentially, $O(N\log N)$ cost at each call site), we still need to deal with large complexity from constructing the linear operators $A$ (i.e., $O(K^{3})$ cost to invert $N+1$ matrices) and performing matrix products with a truncation of $DF(\bar{\chi})$ to estimate $Z_{1}$. The last cost is exacerbated by the use of interval arithmetic, with sufficiently tight enclosures to avoid loss of accuracy. For the proof of stability, naturally, these computational considerations must likewise be taken into account.
Numerical evidence suggests that this instability arises from the occurrence of period-doubling bifurcations along the branch; see Figure 3(b). It may be the case that further decreasing $a_{1}$ results in a cascade of period-doubling bifurcations, eventually leading to a chaotic attractor in $\mathbb{R}_{+}^{3}$ around $a_{1}\approx 4$; see Figure 3(c). If there is indeed a chaotic attractor, its existence, and the mechanism behind its birth deserve further study.
While this article focused on a single Lotka–Volterra model (1), the CAP-based continuation framework we have presented is not limited to this example. In the future, it can be adapted and automated to explore broader families of Lotka–Volterra systems (5) with varying dimensions and interaction structures. This opens the possibility of systematically investigating mechanisms of stable coexistence and exclusion across models that were previously intractable. By providing a rigorous computational approach, our work lays a foundation for a more comprehensive understanding of resource-competition dynamics in complex ecological systems.
@@
Figure 3: The parameter values are $a_{2}=41$, $d_{1}=0.8$, $d_{2}=0.5$, and $m_{1}=m_{2}=y_{1}=y_{2}=\gamma=1$. (a) Numerical approximation of the global family for $a_{1}=6$. The orange rings corresponds to a numerical observation of a period-doubling bifurcation. (b) Numerical approximation of the Floquet multipliers associated to (a). The orange cross marks the crossing through 1-1, i.e., a potential period-doubling bifurcation. (c) Projection onto the $(X_{1},X_{2})$-plane of a numerical chaotic attractor in $\mathbb{R}^{3}_{+}$ for $a_{1}=4$.
@@ 5 Acknowledgments
We are grateful to Sze-Bi Hsu for the suggestion of solving the stable connection problem and to Bernold Fiedler for many inspiring discussions. J.-Y. Dai was supported by NSTC (National Science and Technology Council) under grant No. 113-2628-M-007-005-MY4. O. Hénot was supported by ANR under project No. CAPPS: ANR-23-CE40-0004-01, and by NSTC under grant No. 115-2115-M-002-001-MY2. O. Hénot also thanks the National Center for Theoretical Sciences for its support during the preparation of this work. P. Lappicy was supported by Marie Skłodowska–Curie Actions, UNA4CAREER H2020 Cofund, 847635, with the project DYNCOSMOS. N. Vassena was supported by the DFG (German Research Society), project No. 512355535.
@@ Appendix A Contraction argument for the Floquet normal form
In the context of this manuscript, we want to verify Theorem 2.1 in the Banach space $\mathcal{Y}_{\nu}$ and the fixed-point operator
@@
thereby obtaining a zero of the mapping $G$ defined in (81). Similarly to the construction of the injective linear operator $A$, we set $B\in\mathscr{B}(\mathcal{Y}_{\nu},\mathcal{Y}_{\nu})$ as
$B\coloneqq B_{\textnormal{finite}}\Pi_{K}+B_{\textnormal{tail}}\Pi_{>K},$ (84)
@@
$B_{\textnormal{tail}}\upsilon\coloneqq(0,I_{3\times 3}\otimes V^{\prime}),\qquad(V_{ij}^{\prime})_{k}\coloneqq\begin{cases}0,&|k|\leq K,\\ (ik)^{-1}(V_{ij})_{k},&|k|>K,\end{cases}\quad i,j=1,2,3.$ (85)
@@
[…]
where $C^{T}$ is the 3-by-3 matrix transpose of $C$ and $E_{0}:\textnormal{Mat}_{3\times 3}(\mathcal{U}_{\nu})\to\textnormal{Mat}_{3\times 3}(\mathcal{P}_{\nu})$ is the evaluation operator at 0, i.e. $E_{0}V\coloneqq V(0)$.
We apply Theorem 2.1. Formulas for $\|BG(\bar{\upsilon})\|_{\mathcal{Y}_{\nu}}$, $\|BDG(\bar{\upsilon})-I\|_{\mathscr{B}(\mathcal{Y}_{\nu},\mathcal{Y}_{\nu})}$, and $\sup_{\upsilon\in B_{R}(\bar{\upsilon})}\|BD^{2}G(\upsilon)\|_{\mathscr{B}(\mathcal{Y}_{\nu},\mathcal{Y}_{\nu})}$ are in fact much simpler than those obtained for $F$ in Sections 2.2.12.2.3. Indeed, $G$ has only the single quadratic term $VC$. Thus, we can freely pick $R=\infty$, and $\|BD^{2}G(\upsilon)\|_{\mathscr{B}(\mathcal{Y}_{\nu},\mathcal{Y}_{\nu})}\leq 2\|B\|_{\mathscr{B}(\mathcal{Y}_{\nu},\mathcal{Y}_{\nu})}$ for all $\upsilon\in\mathcal{Y}_{\nu}$.
@@
$\hat{W}_{1}\coloneqq\begin{pmatrix}0&I_{3\times 3}\otimes E_{0}\Pi_{N,K}\\ I_{3\times 3}\otimes\bar{V}&I_{3\times 3}\otimes\partial_{t}+\bar{C}^{T}\otimes I_{3\times 3}-\omega_{1}\end{pmatrix}$, (86)